# Scaling Laws of RoPE-based Extrapolation

### v1

@@ Appendix E Text Continuation
Finally, we execute a text continuation experiment. For LLaMA2 7B(Touvron et al., 2023b) fine-tuned with various base values and fine-tuning lengths, we provide a context of 32K tokens from Books3(Presser, 2020) and prompt the model to continue writing. The resulting texts, after cleaning the special characters, are shown in Figure 12. Remarkably, whether the base is set at 500 or 1000000, the generated text remains coherent, grammatically accurate, and logically consistent. For base 500, extended tuning length or incorporating log-scaled attention Su (2023b) in the testing phase yields improved continuation results. Given the strict extrapolation upper bound for base 1000000, there is a compelling case that a model fine-tuned with base 500 possesses an infinite extrapolation potential.
Context: THIS BOOK IS FOR INFORMATIONAL PURPOSES ONLY. IT IS NOT INTENDED TO BE A SUBSTITUTE FOR CONSULTATION WITH A PROFESSIONAL HEALTHCARE PRACTITIONER. …… They invited airplane pilots with an average age of fifty-two who were interested in participating and evaluated performance in flight simulations. Every three minutes, the pilots were peppered with complex commands and codes that had to be remembered and implemented while under stressful flying conditions. After seven simulations, the study subjects were given Aricept for
Reference: a month and then retested under equally demanding but new simulations. The results showed that they performed better in the later round of evaluations, and the authors stated that Aricept ”appears to have beneficial effects on retention of training on complex aviation tasks in nondemented older adults,” and these findings were ultimately published in Neurology. There is already a major pharmacologic market for physical performance-enhancing drugs. With an aging population, interest in medications for cognitive enhancement will likely grow.
Output (LLaMA2 7B): OOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOO
Output (base=500): the purpose of improving their cognitive performance. Neuro-Geek: Alzheimer’s Drugs for the Healthy? While most adults will not develop Alzheimer’s, many adults will have mild cognitive impairments that affect how they perform in daily life and work. In the sector of the scientific community focused on developing drugs to improve the symptoms associated with dementia, there is an understanding that those same drugs may be used by healthy people with the goal of maintaining cognitive acuity or even, in certain circumstances, improving mental agility above their established baseline.
Output (base=1000000): a month. The results were clear: the drug improved their performance. The study was published in 2015, and the authors concluded that ”the results suggest that the drug may be able to improve cognitive performance in healthy individuals.” The study was small, and the results are not yet conclusive. But it is a start. Prescription Antidepressants The first antidepressants were developed in the 1950s, and they have been used to treat depression ever since. But in the 1990s, scientists began to wonder if they might also be useful for improving cognitive functioning.
Output (base=500 16K): a week. The drug is a cholinesterase inhibitor, which means it blocks the breakdown of acetylcholine, a neurotransmitter that helps regulate memory and learning. The results were striking. The pilots who took the drug performed better than those who didn’t. They were able to remember more of the complex commands and codes, and they were able to implement them more quickly. The study authors concluded that the drug ”may be useful in improving cognitive performance in healthy individuals.” But there’s a catch. The drug is not approved for use in healthy people. It’s only approved for people with Alzheimer’s disease.
Output (base=1000000 16K): a month, then switched to a placebo for another month. The results were striking. The pilots who took the drug improved their performance by 20 percent, while those who took the placebo got worse. The study was small, and the results were not replicated in a larger study. But the idea that healthy people might be able to improve their cognitive performance with drugs designed to treat dementia is intriguing. Prescription Antidepressants The most widely prescribed class of antidepressants, selective serotonin reuptake inhibitors (SSRIs), are also used to treat anxiety, obsessive-compulsive disorder, and panic disorder.
Output (base=500 16K log-scaled): a week. The drug is a cholinesterase inhibitor, which means it blocks the breakdown of acetylcholine, a neurotransmitter that is critical for memory and learning. The results were striking. The pilots who took the drug performed better than those who didn’t. They were able to remember more of the complex commands and codes and to implement them more quickly. But the study also had a major flaw. The pilots were not randomly assigned to take the drug or not. They were all given the drug because they were already healthy. The study’s authors concluded that the drug improved performance in pilots who were already healthy. But that’s not what the study actually showed.
@@
Limitation
In this work, our primary objective is to elucidate the mechanisms by which RoPE, using either smaller or larger bases, enhances extrapolation performance. Furthermore, there remains a need to rigorously assess our method on extended context data. Additionally, it needs additional efforts to integrate whether there exist upper and lower bounds for adjusting the base. Finally, since the base value and pre-training context length is the key to extrapolation, there might be an optimal pair of base and training length for RoPE-based LLM pre-training to be discovered.

### v2

@@ Appendix E Text Continuation
Finally, we execute a text continuation as our case study. For LLaMA2 7B(Touvron et al., 2023b) fine-tuned with various base values and fine-tuning lengths, we provide a context of 32K tokens from Books3(Presser, 2020) and prompt the model to continue writing. The resulting texts, after cleaning the special characters, are shown in Figure 12. Remarkably, whether the base is set at 500 or 1000000, the generated text remains coherent, grammatically accurate, and logically consistent. For base 500, extended tuning length or incorporating log-scaled attention Su (2023b) in the testing phase yields improved continuation results. Given the strict extrapolation upper bound for base 1000000, there is a compelling case that a model fine-tuned with base 500 possesses an infinite extrapolation potential.
Limitation
In this work, our primary objective is to elucidate the mechanisms by which RoPE, using either smaller or larger bases, enhances extrapolation performance. Additionally, there remains a need to find whether there exist upper and lower bounds for adjusting the base. Besides, it needs additional efforts to integrate the performance of RoPE with smaller bases in the pre-training stage. Finally, since the base value and pre-training context length are keys to extrapolation, there might be an optimal pair of base and training length for RoPE-based LLM pre-training to be discovered.
Context: THIS BOOK IS FOR INFORMATIONAL PURPOSES ONLY. IT IS NOT INTENDED TO BE A SUBSTITUTE FOR CONSULTATION WITH A PROFESSIONAL HEALTHCARE PRACTITIONER. …… They invited airplane pilots with an average age of fifty-two who were interested in participating and evaluated performance in flight simulations. Every three minutes, the pilots were peppered with complex commands and codes that had to be remembered and implemented while under stressful flying conditions. After seven simulations, the study subjects were given Aricept for Reference: a month and then retested under equally demanding but new simulations. The results showed that they performed better in the later round of evaluations, and the authors stated that Aricept ”appears to have beneficial effects on retention of training on complex aviation tasks in nondemented older adults,” and these findings were ultimately published in Neurology. There is already a major pharmacologic market for physical performance-enhancing drugs. With an aging population, interest in medications for cognitive enhancement will likely grow. Output (LLaMA2 7B): OOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOO Output (base=500): the purpose of improving their cognitive performance. Neuro-Geek: Alzheimer’s Drugs for the Healthy? While most adults will not develop Alzheimer’s, many adults will have mild cognitive impairments that affect how they perform in daily life and work. In the sector of the scientific community focused on developing drugs to improve the symptoms associated with dementia, there is an understanding that those same drugs may be used by healthy people with the goal of maintaining cognitive acuity or even, in certain circumstances, improving mental agility above their established baseline. Output (base=1000000): a month. The results were clear: the drug improved their performance. The study was published in 2015, and the authors concluded that ”the results suggest that the drug may be able to improve cognitive performance in healthy individuals.” The study was small, and the results are not yet conclusive. But it is a start. Prescription Antidepressants The first antidepressants were developed in the 1950s, and they have been used to treat depression ever since. But in the 1990s, scientists began to wonder if they might also be useful for improving cognitive functioning. Output (base=500 16K): a week. The drug is a cholinesterase inhibitor, which means it blocks the breakdown of acetylcholine, a neurotransmitter that helps regulate memory and learning. The results were striking. The pilots who took the drug performed better than those who didn’t. They were able to remember more of the complex commands and codes, and they were able to implement them more quickly. The study authors concluded that the drug ”may be useful in improving cognitive performance in healthy individuals.” But there’s a catch. The drug is not approved for use in healthy people. It’s only approved for people with Alzheimer’s disease. Output (base=1000000 16K): a month, then switched to a placebo for another month. The results were striking. The pilots who took the drug improved their performance by 20 percent, while those who took the placebo got worse. The study was small, and the results were not replicated in a larger study. But the idea that healthy people might be able to improve their cognitive performance with drugs designed to treat dementia is intriguing. Prescription Antidepressants The most widely prescribed class of antidepressants, selective serotonin reuptake inhibitors (SSRIs), are also used to treat anxiety, obsessive-compulsive disorder, and panic disorder. Output (base=500 16K log-scaled): a week. The drug is a cholinesterase inhibitor, which means it blocks the breakdown of acetylcholine, a neurotransmitter that is critical for memory and learning. The results were striking. The pilots who took the drug performed better than those who didn’t. They were able to remember more of the complex commands and codes and to implement them more quickly. But the study also had a major flaw. The pilots were not randomly assigned to take the drug or not. They were all given the drug because they were already healthy. The study’s authors concluded that the drug improved performance in pilots who were already healthy. But that’s not what the study actually showed.